# The T2K Collaboration Signal selection and model-independent extraction of the neutrino neutral-current single cross section with the T2K experiment K. Abe

### v2

@@
††thanks: also at Université Paris-Saclay††thanks: also at Departament de Fisica de la Universitat Autonoma de Barcelona, Barcelona, Spain.††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at Moscow Institute of Physics and Technology (MIPT), Moscow region, Russia and National Research Nuclear University ”MEPhI”, Moscow, Russia††thanks: also at IPSA-DRII, France††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at JINR, Dubna, Russia††thanks: also at the Graduate University of Science and Technology, Vietnam Academy of Science and Technology††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at Nambu Yoichiro Institute of Theoretical and Experimental Physics (NITEP)††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at Nambu Yoichiro Institute of Theoretical and Experimental Physics (NITEP)††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan
@@ Abstract
This article presents a study of single production in neutrino neutral-current interactions (NC1) using the ND280 detector of the T2K experiment. We report the largest sample of such events selected by any experiment, providing the first new data for this channel in over four decades and the first using a sub-GeV neutrino flux. The signal selection strategy and its performance are detailed together with validations of a robust cross section extraction methodology. The measured flux-averaged integrated cross-section is , 1.3 above the NEUT v5.4.0 expectation.
@@ III Event Simulation and Signal Selection
Proton beam interactions with the graphite target and the propagation of daughter particles through the target station are modeled using FLUKA Ferrari et al. (2005); Böhlen et al. (2014). Subsequent interactions and decays in the beamline are modeled with GEANT3 Brun et al. (1994) and GCALOR Zeitnitz and Gabriel (1992) to predict the neutrino flux. The run-by-run flux simulation is tuned to the beam conditions as recorded by the beam monitoring systems and hadronic interactions and multiplicities are tuned to the NA61/SHINE hadron production measurements with a T2K replica target Abgrall et al. (2016, 2019). Neutrino interactions in the ND280 detector including the description of all the final-state particle kinematics are simulated using the NEUT v5.4.0 Hayato (2009); Hayato and Pickering (2021) and GENIE v2.8.0 Andreopoulos et al. (2010) neutrino event generators, hereafter NEUT and GENIE. An in-depth description on the details for the NEUT (GENIE) model can be found in Ref. Abe et al. (2023a) (Ref. Abe et al. (2020a)). The propagation of the final-state particles through the detector medium is simulated using GEANT4 v4.9.4 Agostinelli et al. (2003); Allison et al. (2006, 2016) with the QGSP_BERT physics list Allison et al. (2016). Pion secondary interactions are handled by the cascade model in NEUT.
@@ III.1 Signal and Background Definitions
Simulated events are classified according to the following definitions, based on their true information after Final-State Interactions (FSI), i.e. considering only particles that exit the nucleus:
CC0: interactions with a muon and without mesons in the final state.
CC1: interactions with a muon and a positive pion and no other mesons in the final state.
CC-other: interactions with a muon which are not included in other topologies.
CC: interactions with an antimuon in the final state.
/ CC: and interactions with an electron or positron in the final state.
NC0: and interactions without charged leptons nor mesons in the final state.
NC1 0p (Signal): and interactions without charged leptons, a single positive pion, no protons with true momentum MeV/ and no other mesons nor other charged particles in the final state. The cut on true proton momentum is discussed in the following section (Sec. III.3).
@@
NCX: and interactions without charged leptons and at least one neutral pion in the final state. Any number of other mesons are allowed.
NC-other: and interactions without a charged lepton that are not included in any other topology.
@@ III.2 Signal Selection Method
Due to its low cross section, NC1 events are expected to make up only 1.3% of all interactions in FGD1. A selection algorithm was developed to identify a subset of events enriched in signal, contained in a signal sample, complemented by background-enriched samples. The logic of the selection algorithm, associated with Figure 3, is as follows:
@@
FGD1-TPC: The main track in the event, i.e. the only track that exits FGD1, must reach at least one ND280 TPC, creating a reconstructed track segment of sufficient length to ensure good TPC information. Events with multiple tracks that escape the FGD1 volume are rejected, as for signal events only one track is expected. Events with one additional track contained in FGD1 are retained separately, in a dedicated sample, that will be referred to as the Additional Track Sample (AddTrkS).
@@
TPC PID: The TPC particle identification (PID) information —based on the measured with respect to the particle momentum measured by curvature— must be compatible with that of a Minimum Ionizing Particle (MIP). Tracks rejected by this cut are retained separately in the TPC PID Sample (TPIDS).
@@
ECal PID: For those tracks reaching the ECal, their PID information must be compatible with a hadron. The ECal offers excellent capabilities in separating vs charged , and in consequence, rejected events are collected in a separate control sample, named the ECal PID Sample (EPIDS).
@@
Figure 3 shows the efficiency and purity of the event selection for events with 0.5 and GeV/ (see Sec.III.4). The main efficiency drops correspond to the necessary requirement of having the candidate reaching the TPC and to the ECal PID cut. Notably, events rejected by the ECal PID are largely retained in the EPIDS. The purity, negligible before applying PID cuts, increases drastically with the last cuts in the selection, demonstrating an effective enhancement of the fraction of signal events with a minor decrease in efficiency. The impact of the cuts on the signal efficiency and purity are well-matched in NEUT and GENIE simulated events demonstrating that the selection process is robust to the simulation physics model.
@@ III.3 Proton Threshold for Signal Events
NC1 events with proton momentum below 200 MeV/ are sometimes selected and are therefore included in the signal definition to be robust against model prediction discrepancies. Events with protons above 200 MeV/ are rarely selected, with much lower selection efficiency than other NC1 events. Therefore, they are treated as background. To validate this choice, fake data studies modifying the number of background NC1 events with protons were performed, and as later reported in Sec. V, no measurement bias was observed.
@@ III.4 Kinematic Considerations About The Positive Pion
Figure 5: Projected 1D efficiencies as a function of the true momentum and angle. The top two panels are for the full phase space and the bottom two for phase space restricted to the RoI, see the text for details.
@@ III.5 Signal Sample
Table 1 shows the total expected events for NEUT and GENIE. The expected signal events double the largest previous existing sample, provided by the Gargamelle Krenz et al. (1978). An additional 56.3 (53.1) signal events are selected in the EPIDS according to NEUT (GENIE), with an overall purity similar to 10%, reaching 30% at around 300 MeV/. In total, this analysis selects over 200 NC events, most of them in regions with significant purity.
@@ III.6 Background Samples
Figure 7: One dimensional momentum and angle distributions for the background-enriched sample. From top to bottom the plots in each row corresponds to the EPIDS, the AddTrkS and the TPIDS.
@@
EPIDS: Consists of signal-like events where the main track has an ECal PID consistent with that of a instead of a .
AddTrkS: Consists of signal-like events with an additional reconstructed track contained in FGD1.
TPIDS: Consists of signal-like events where the main track is not compatible with a according to TPC PID information.
The expected and observed event rates in all samples are summarized in Table 2. In all samples GENIE, NEUT and data follow similar distributions both in the selected track reconstructed momentum and its angle. A deficit of data events is observed in the forward-most bin in the EPIDS, a feature that has been observed in the past in CC0 studies, e.g. Refs. Rodrigues et al. (2016); Ruterbories et al. (2019); Abe et al. (2020b, c), and that is attributed to a mis-modeling of neutrino interactions with low squared momentum transfer, . As this is a known feature, we tested the robustness of the signal cross section extraction to modifications and no bias was observed. Further details are later presented in Sec. V.
Table 2: Total number of expected selected events in the Signal Region (SR) and in the EPIDS, AddTrkS and TPIDS for NEUT and GENIE normalized to the data POT
@@ IV.1 Binned Likelihood Fit
The signal and control samples are binned in reconstructed track momentum and angle. The expected event rate per bin is calculated by varying the nominal predictions according to model uncertainties. These variations account for uncertainties in the flux, detector and cross-section parameters. Additionally, the NC1 signal cross section is adjusted using one free parameter per bin, referred to as template parameters. The binning scheme, detailed in Appendix A, consists of 13 (85) bins in true (reconstructed) track angle and momentum. The true kinematic bins, where the cross section is extracted, are shown in Fig. 8. The parameters are simultaneously optimized to maximize agreement with data in both signal and background samples. This optimization provides a data-driven background constraint while unfolding detector effects and estimating the number of signal events in each bin.
@@ IV.2.1 Flux Model Uncertainty
All events are grouped in bins of flavor (, , , ) and true neutrino energy (20 bins per flavor). There is one fit parameter per flux bin, for a total of 80 flux parameters. These apply identical scaling weights to all associated events. Uncertainties across all flux bins are correlated and characterized using a flux covariance matrix produced by simultaneously varying underlying flux model parameters in the simulation according to their prior uncertainties. For the flavor, contributing to the vast majority of events, the uncertainty is about 5% Vladisavljevic (2018) for the most important kinematic neutrino energy region, i.e. around the flux peak at E 0.6 GeV.
@@ IV.2.2 Detector Model Uncertainty
One parameter is associated to every reconstructed kinematic bin, for a total of 85 bins summarized in Tables 5, 6, 7 and 8 in Appendix A. Each detector parameter consists of a normalization weight that equally scales all signal and background events in that reconstructed bin. The prior error on each detector parameter is characterized by a detector covariance matrix calculated for this study using an ensemble of 500 selected event rate distributions that are obtained varying simultaneously all detector effects based on their prior uncertainties (e.g. PID variables, reconstruction efficiencies, amount of pile-up, etc). For the bins in the signal sample the integrated rate uncertainty in the RoI is 4.5%, dominated by the modeling of pion secondary interactions (SI) (2.0%), the ECal PID (1.9%), and the TPC PID (1.9%). The overall level of uncertainty is also below 5% for the EPIDS and TPIDS, with the ECal PID (TPC PID) being the dominant uncertainty for the former (latter) at the level of 2.8% (3.0%). For the AddTrkS the integrated uncertainty is of 8.6%, dominated by the modeling of pion SI (3.7%) and the main track momentum resolution (3.5%).
@@ IV.2.3 Cross-Section Model Uncertainty
A total of 28 cross section model parameters are considered. Unless otherwise specified their values and uncertainties are those in Ref. Abe et al. (2023b).
@@
Six FSI parameters related to inelastic interactions, pion absorption and charge exchange below and above 0.5 GeV.

### v3

@@
††thanks: also at Université Paris-Saclay††thanks: also at Departament de Fisica de la Universitat Autonoma de Barcelona, Barcelona, Spain††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at Moscow Institute of Physics and Technology (MIPT), Moscow region, Russia and National Research Nuclear University ”MEPhI”, Moscow, Russia††thanks: also at IPSA-DRII, France††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at JINR, Dubna, Russia††thanks: also at the Graduate University of Science and Technology, Vietnam Academy of Science and Technology††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at Nambu Yoichiro Institute of Theoretical and Experimental Physics (NITEP)††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at Nambu Yoichiro Institute of Theoretical and Experimental Physics (NITEP)††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan
@@ A. Bubak
J.A. Caballero
Departamento de Física Atómica, Molecular y Nuclear, Universidad de Sevilla, 41080 Sevilla, Spain
F. Cadoux
@@ C. Forza
J. García-Marcos
Ghent University, Department of Physics and Astronomy, Proeftuinstraat 86, B-9000 Gent, Belgium
@@ V. Glagolev
R. González Jiménez
Departamento de Física Atómica, Molecular y Nuclear, Universidad de Sevilla, 41080 Sevilla, Spain
J. González Rosa
Departamento de Física Atómica, Molecular y Nuclear, Universidad de Sevilla, 41080 Sevilla, Spain
@@ A. Holin
M.C.F. Hooft
Ghent University, Department of Physics and Astronomy, Proeftuinstraat 86, B-9000 Gent, Belgium
@@ A. Izmaylov
N. Jachowicz
Ghent University, Department of Physics and Astronomy, Proeftuinstraat 86, B-9000 Gent, Belgium
@@ A. Konaka
A. Kumar Jha
Ghent University, Department of Physics and Astronomy, Proeftuinstraat 86, B-9000 Gent, Belgium
@@ A. Mefodiev
G.D. Megias
Departamento de Física Atómica, Molecular y Nuclear, Universidad de Sevilla, 41080 Sevilla, Spain
@@ V.Q. Nguyen
K. Niewczas
Ghent University, Department of Physics and Astronomy, Proeftuinstraat 86, B-9000 Gent, Belgium
@@ Abstract
This article presents a study of single production in neutrino neutral-current interactions (NC1) using the FGD1 hydrocarbon target of the ND280 detector of the T2K experiment. We report the largest sample of such events selected by any experiment, providing the first new data for this channel in over four decades and the first using a sub-GeV neutrino flux. The signal selection strategy and its performance are detailed together with validations of a robust cross section extraction methodology. The measured flux-averaged integrated cross-section is , 1.3 above the NEUT v5.4.0 expectation.
@@ III Event Simulation and Signal Selection
Proton beam interactions with the graphite target and the propagation of daughter particles through the target station are modeled using FLUKA Ferrari et al. (2005); Böhlen et al. (2014). Subsequent interactions and decays in the beamline are modeled with GEANT3 Brun et al. (1994) and GCALOR Zeitnitz and Gabriel (1992) to predict the neutrino flux. The run-by-run flux simulation is tuned to the beam conditions as recorded by the beam monitoring systems and hadronic interactions and multiplicities are tuned to the NA61/SHINE hadron production measurements with a T2K replica target Abgrall et al. (2016, 2019). Neutrino interactions in the ND280 detector including the description of all the final-state particle kinematics are simulated using the NEUT v5.4.0 Hayato (2009); Hayato and Pickering (2021) and GENIE v2.8.0 Andreopoulos et al. (2010) neutrino event generators, hereafter NEUT and GENIE, using respectively about 20 (10) times more POT than the data. An in-depth description on the details for the NEUT (GENIE) model can be found in Ref. Abe et al. (2023a) (Ref. Abe et al. (2020a)). The propagation of the final-state particles through the detector medium is simulated using GEANT4 v4.9.4 Agostinelli et al. (2003); Allison et al. (2006, 2016) with the QGSP_BERT physics list Allison et al. (2016). Pion secondary interactions are handled by the cascade model in NEUT and treated as a detector systematic uncertainty.
@@ III.1 Signal and Background Definitions
Simulated events are classified according to the following definitions, based on their true information after Final-State Interactions (FSI), i.e. considering only particles that exit the nucleus. The categories are built based on the particle content irrespectively of its kinematic except for the NC1 0p and NC1 Np categories are descried below:
CC0: All interactions with a muon and without mesons in the final state.
CC1: All interactions with a muon and a positive pion and no other mesons in the final state.
CC-other: All interactions with a muon which are not included in other topologies.
CC: All interactions with an antimuon in the final state.
/ CC: All and interactions with an electron or positron in the final state.
NC0: All and interactions without charged leptons nor mesons in the final state.
NC1 0p (Signal): All and interactions without charged leptons, a single positive pion, no protons with true momentum MeV/ and no other mesons nor other charged particles in the final state. Any number of neutrons is allowed. The cut on true proton momentum is discussed in the following section (Sec. III.3).
@@
NCX: All and interactions without charged leptons and at least one neutral pion in the final state. Any number of other mesons are allowed.
NC-other: All and interactions without a charged lepton that are not included in any other topology.
@@ III.2 Signal Selection Method
Due to its low cross section, NC1 events are expected to make up only 1.3% of all interactions in FGD1. A selection algorithm was developed to identify a signal sample enriched in signal events, complemented by background-enriched samples. The logic of the selection algorithm, associated with Figure 3, is as follows:
@@
FGD1-TPC: The main track in the event, i.e. the only track that exits FGD1, must reach at least one ND280 TPC, creating a reconstructed track segment. This TPC segment must have at least 19 hit clusters, a standard criteria in analyses, to ensure reliable TPC reconstructed information. Events that consist exclusively of FGD1-contained tracks are rejected. Events with multiple tracks that escape the FGD1 volume are rejected, as for signal events only one track is expected. Events with one additional track contained in FGD1 are retained separately, in a dedicated sample, that will be referred to as the Additional Track (AddTrk) sample.
@@
TPC PID: The TPC particle identification (PID) information —based on the measured with respect to the particle momentum measured by curvature— must be compatible with that of a Minimum Ionizing Particle (MIP). Tracks rejected by this cut are retained separately in the TPC PID (TPID) sample.
@@
ECal PID: For those tracks reaching the ECal, their PID information must be compatible with a hadron. The ECal offers excellent capabilities in separating vs charged , and in consequence, rejected events are collected in a separate control sample, named the ECal PID (EPID) sample.
@@
Figure 3 shows the efficiency and purity of the event selection for events with 0.5 and GeV/ (see Sec.III.4). The main efficiency drops correspond to the necessary requirement of having the candidate reaching the TPC and to the ECal PID cut. Notably, events rejected by the ECal PID are largely retained in the EPID sample. The purity, negligible before applying PID cuts, increases drastically with the last cuts in the selection, demonstrating an effective enhancement of the fraction of signal events with a minor decrease in efficiency. The impact of the cuts on the signal efficiency and purity are well-matched in NEUT and GENIE simulated events demonstrating that the selection process is robust to the simulation physics model.
@@ III.3 Proton Threshold for Signal Events
NC1 events with proton momentum below 200 MeV/ are sometimes selected and are therefore included in the signal definition to be robust against model prediction discrepancies. Events with protons above 200 MeV/ are rarely selected, with much lower selection efficiency than other NC1 events. Therefore, they are treated as background. NEUT, modeling FSI using a semi-classical intra-nuclear cascade model, and GENIE using the “hA” model, render noticeably differences as observed Fig. 4. To validate the proton threshold choice, fake data studies modifying the number of background NC1 events with protons were performed. As later reported in Sec. V, no measurement bias was observed.
@@ III.4 Kinematic Considerations About The Positive Pion
Figure 5: Projected 1D efficiencies as a function of the true momentum and angle. The top two panels are for the full phase space and the bottom two for phase space restricted to the RoI, see the text for details. Error bars show the MC statistical uncertainty.
@@ III.5 Signal Sample
Table 1 shows the total expected events for NEUT and GENIE. The expected signal events double the largest previous existing sample, provided by the Gargamelle Krenz et al. (1978). An additional 56.3 (53.1) signal events are selected in the EPID sample according to NEUT (GENIE), with an overall purity similar to 10%, reaching 30% at around 300 MeV/. In total, this analysis selects over 200 NC events, most of them in regions with significant purity.
@@ III.6 Background Samples
Figure 7: One dimensional momentum and angle distributions for the background-enriched sample. From top to bottom the plots in each row corresponds to the EPID, the AddTrk and the TPID sample.
@@
EPID: Consists of signal-like events where the main track has an ECal PID consistent with that of a instead of a .
AddTrk: Consists of signal-like events with an additional reconstructed track contained in FGD1.
TPID: Consists of signal-like events where the main track is not compatible with a according to TPC PID information.
The expected and observed event rates in all samples are summarized in Table 2. In all samples GENIE, NEUT and data follow similar distributions both in the selected track reconstructed momentum and its angle. A deficit of data events is observed in the forward-most bin in the EPID sample, a feature that has been observed in the past in CC0 studies, e.g. Refs. Rodrigues et al. (2016); Ruterbories et al. (2019); Abe et al. (2020b, c), and that is attributed to a mis-modeling of neutrino interactions with low squared momentum transfer, . As this is a known feature, we tested the robustness of the signal cross section extraction to modifications and no bias was observed. Further details are later presented in Sec. V.
Table 2: Total number of expected selected events in the Signal Region (SR) and in the EPID, AddTrk and TPID samples for NEUT and GENIE normalized to the data POT
@@ IV.1 Binned Likelihood Fit
The signal and control samples are binned in reconstructed track momentum and angle. The expected event rate per bin is calculated by varying the nominal predictions according to model uncertainties. These variations account for uncertainties in the flux, detector and cross-section parameters. Additionally, the NC1 signal cross section is adjusted using one free parameter per bin, referred to as template parameters. The binning scheme, detailed in Appendix B, consists of 13 (85) bins in true (reconstructed) track angle and momentum. The true kinematic bins, where the cross section is extracted, are shown in Fig. 8. The parameters are simultaneously optimized to maximize agreement with data in both signal and background samples. This optimization provides a data-driven background constraint while unfolding detector effects and estimating the number of signal events in each bin.
@@ IV.2.1 Flux Model Uncertainty
All events are grouped in bins of flavor (, , , ) and true neutrino energy (20 bins per flavor). There is one fit parameter per flux bin, for a total of 80 flux parameters. These apply identical scaling weights to all associated events. Uncertainties across all flux bins are correlated and characterized using a flux covariance matrix produced by simultaneously varying underlying flux model parameters in the simulation according to their prior uncertainties. For the flavor prediction, contributing to the vast majority of events, the uncertainty is about 5% Vladisavljevic (2018) for the most important kinematic neutrino energy region, i.e. around the flux peak at E 0.6 GeV.
@@ IV.2.2 Detector Model Uncertainty
One parameter is associated to every reconstructed kinematic bin, for a total of 85 bins summarized in Tables 5, 6, 7 and 8 in Appendix B. Each detector parameter consists of a normalization weight that equally scales all signal and background events in that reconstructed bin. The prior error on each detector parameter is characterized by a detector covariance matrix calculated for this study using an ensemble of 500 distributions of the events selected obtained by varying simultaneously all detector effects based on their prior uncertainties (e.g. PID variables, reconstruction efficiencies, pion and proton secondary interactions, amount of pile-up, etc). Therefore, the final event rate variations reflect changes in the selection efficiency due to changes in variables used to define the selection criteria, and account for the migration of the selected track kinematics. For the bins in the signal sample the integrated rate uncertainty in the RoI is 4.5%, dominated by the modeling of pion secondary interactions (SI) (2.0%), the ECal PID (1.9%), and the TPC PID (1.9%). The overall level of uncertainty is also below 5% for the EPID and TPID samples, with the ECal PID (TPC PID) being the dominant uncertainty for the former (latter) at the level of 2.8% (3.0%). For the AddTrk sample the integrated uncertainty is of 8.6%, dominated by the modeling of pion SI (3.7%) and the main track momentum resolution (3.5%).
@@ IV.2.3 Cross-Section Model Uncertainty
A total of 28 cross section model parameters are considered. Unless otherwise specified their prior uncertainties and modeling match the description in Ref. Abe et al. (2023b).
@@
Six FSI parameters related to inelastic interactions, pion absorption, production and charge exchange below and above 0.5 GeV.